# Towards deployment-centric multimodal AI beyond vision and language

### v1

@@ Introduction
Deploying multimodal AI systems in pandemic response translates research into actionable solutions for high-stakes environments. For instance, patient surge management can integrate biosignals, medical images, and clinical notes to optimise intensive care unit bed allocation and staff deployment. Public health decision-making can benefit from multimodal analyses of epidemiological data, EHRs, and social media trends to guide interventions such as lockdowns or resource distribution. Real-time deployment requires robust infrastructure to process asynchronous data streams and address modality-specific interruptions, such as gaps in biosignal or social media data. Adaptive algorithms and reliable systems ensure timely and accurate outbreak prediction and monitoring. Assessment through multi-centre trials and multi-disciplinary benchmarks mincu2022developing can validate the added value of multimodal AI over unimodal systems. Measuring the accuracy of outbreak prediction or the effectiveness of intervention informed by multimodal AI can ensure trust and accountability. Further addressing scalability and privacy compliance can support sustainable and ethical deployment.
@@
Self-driving car design (Supplementary Fig. S3) exemplifies a transformative application of multimodal AI, addressing challenges in safety, efficiency, and sustainability soni2021design; badue2021self. Defining the problem involves determining how multimodal AI can significantly outperform unimodal approaches for navigating complex environments safely and efficiently. For instance, fusing sensors such as LiDAR (light detection and ranging), radar and cameras can enhance perception and decision-making in dynamic settings, especially where a single sensor may fail yeong2021sensor. Beyond perception, multimodal AI can support physical simulations, such as combining text, images, and videos, to model vehicle dynamics and road interactions yang2024generalized. In materials design, integrating structural and chemical data aids in developing lightweight, sustainable materials for vehicle construction zhang2022advanced. Clearly defining these challenges ensures multimodal AI solutions are effectively tailored to the demands of autonomous driving.
@@
The deployment of multimodal AI in self-driving car design faces safety, privacy, technical and economic constraints that must be addressed for safe and scalable adoption. Privacy concerns arise from integrating multimodal data, such as LiDAR, cameras and passenger-related information, which increases the potential for data misuse or unauthorised tracking hansson2021self. Mitigating these risks requires robust data anonymisation and adherence to privacy regulations chowdhury2020attacks. Technical challenges include real-time data processing in dynamic environments, such as urban intersections or adverse weather. Modality-specific failures, such as LiDAR disruptions in heavy rain, require fallback mechanisms such as radar to maintain reliability. Self-driving systems must achieve high predictive accuracy for collision avoidance and route planning, and operate reliably across regions with varying regulations and infrastructure. Economic constraints involve balancing the cost of high-resolution sensors with scalability needs. Meeting sustainability goals requires energy-efficient operation, with trade-offs between sensor performance and power consumption. Addressing these constraints holistically enables robust and accessible deployment of self-driving technologies.
@@
In self-driving car design, task formulation maps problems and constraints to concrete AI objectives for perception, navigation, and safety. A key task is developing AI driving systems that integrate multimodal data streams, including LiDAR, radar, GPS, and cameras, for real-time navigation. These systems must dynamically process inputs to handle traffic, obstacles, and weather variability. For example, integrating road and weather data can enhance route planning and safety in adverse conditions dey2014potential. Another task is predicting material properties and performance to support vehicle design. By fusing structural, experimental, and supply chain data, multimodal AI can optimise materials for lightweight, durable, and cost-effective vehicles, aligning with sustainability goals kamran2022artificial. Additional tasks include passenger-focused objectives, such as integrating biosignals and cabin sensors to enhance comfort and safety. Ensuring privacy compliance remains integral across all tasks.
@@
The development of multimodal AI systems for self-driving cars follows the five-step workflow outlined above, adapted to the demands of autonomous systems. Data collection involves assembling sensor inputs (e.g. LiDAR, radar, cameras) and materials data to build comprehensive datasets. Data curation addresses challenges such as synchronising asynchronous sensor inputs and standardising materials information. Learning can employ fusion strategies to enhance system reliability, while physics-informed neural networks raissi2019physics; karniadakis2021physics can integrate scientific laws to improve realistic simulations of driving scenarios and materials development. Evaluation ensures robustness through multi-metric validation across diverse scenarios, including urban environments and extreme weather conditions. Finally, interpretation in autonomous systems supports real-time explainability and transparency, enabling stakeholders to trust and audit system decisions. By aligning these steps with the unique requirements of autonomous systems, multimodal AI can ensure adaptability, safety, and performance.
@@
Real-world deployment of multimodal AI systems for self-driving cars requires live sensor integration, infrastructure readiness, and operational resilience across diverse environments. Robust in-vehicle computing and inter-vehicle communication enable reliable and coordinated performance at scale. Assessment validates multimodal AI’s added value through simulations and multi-centre trials. Testing performance under extreme conditions, such as adverse weather or high-traffic intersections, ensures reliability. Addressing scalability and privacy compliance further supports ethical, sustainable deployment, advancing societal trust in self-driving technologies liu2019public.
@@
Climate change adaptation (Supplementary Fig. S3) involves forecasting extreme weather, assessing climate risks, and managing resources to minimise environmental and socio-economic impacts bi_accurate_2023; lam2023learning; mathiesen2018rating; breitenstein2022disclosure. Defining the problem involves assessing whether multimodal AI can surpass unimodal approaches by integrating diverse data sources for deeper insights and actionable strategies. For example, multimodal AI can combine satellite imagery, historical climate records, and geospatial data to model weather patterns and predict extreme events, supporting proactive disaster management imran2020using. Integrating socio-economic data with environmental observations can enable more comprehensive climate risk assessments tanir2024social.
@@
Deploying multimodal AI for climate change adaptation must navigate privacy, technical and economic constraints. Privacy concerns arise from integrating sensitive data, such as socio-economic records and proprietary satellite imagery. Ensuring compliance with global and regional privacy regulations and the ethical use of data supports responsible deployment. Technical challenges include the heterogeneity and sparsity of environmental and socio-economic data. Economic constraints, such as the high cost of advanced sensors, necessitate balancing low-cost and low-fidelity data sources with high-cost, high-fidelity alternatives to maintain accessibility and scalability. By addressing these constraints, multimodal AI can deliver equitable and sustainable climate solutions in diverse global contexts.
@@
For climate change adaptation, task formulation aligns complex environmental challenges with actionable AI tasks. A primary task is forecasting extreme weather events by integrating multimodal data such as satellite imagery, historical climate records, and real-time observations. These forecasts can enable timely interventions to mitigate human and economic losses kumar2021overview, including early warning systems for floods that integrate geospatial, hydrological, and social data streams tkachenko2017predicting. Another task is predicting the socio-economic impacts of severe weather by combining spatial, market, and socio-economic data, guiding resource allocation and policy decisions thulke2024climategpt. These tasks align multimodal AI with both technical requirements and societal priorities, advancing climate resilience and equitable adaptation strategies.
@@
Developing multimodal AI systems for climate change adaptation follows a similar five-step workflow. Data collection gathers remote sensing data, near-real-time sensor network outputs, historical climate records, and socio-economic metrics, creating comprehensive datasets for tasks such as extreme weather forecasting and early warning systems. Data curation addresses challenges such as temporal misalignment and data sparsity by standardising inputs and enriching sparse data through interpolation or integration of auxiliary sources. Multimodal learning can benefit from advanced models such as Aurora bodnar2024aurora, which fuse satellite imagery and meteorological data to improve atmospheric predictions. Evaluation uses multi-metric benchmarks to test system performance in diverse scenarios. For example, WeatherBench 2 provides a standard platform for assessing atmospheric prediction models rasp2024weatherbench. Interpretation ensures outputs are transparent and actionable for stakeholders, supporting evidence-based decision-making. By aligning these steps with climate-specific challenges, multimodal AI can deliver reliable, scalable, and adaptable solutions.
@@
Real-world deployment of multimodal AI for climate change adaptation translates research insights into scalable, operational systems. Early warning platforms and impact forecasting models must synchronise diverse data streams, deliver real-time outputs, and integrate with decision-making infrastructures across sectors and regions morshed2024decoding; Reichstein2025EarlyWO. Real-time monitoring requires robust infrastructure to synchronise and process diverse modalities. Deployment challenges include latency, data coverage gaps, and variable infrastructure capacity across geographies. Assessment involves stress testing and multi-centre validation to ensure reliability across environmental and socio-economic scenarios. Ethical considerations, such as prioritising vulnerable communities and ensuring equitable access to data and AI tools, help enable sustainable deployment. By overcoming deployment challenges, multimodal AI systems can empower policymakers, industries, and communities to respond proactively to climate challenges, advancing global sustainability efforts.
@@
Most existing research on multimodal AI has focused on model-centric development. While attention to data-centric development is increasing, unlocking the full potential of multimodal AI and addressing real-world challenges will ultimately require a shift towards deployment-centric development. This shift will require strategic advancements in data, model, and deployment methods, as well as concerted efforts to address multimodal-AI-specific challenges, namely modality incompleteness, multimodal heterogeneity, cross-modality alignment, modality complementarity, and multimodal privacy risk, through interdisciplinary collaborations and community-building. Box 1 presents strategic recommendations to advance multimodal AI across disciplines under a deployment-centric perspective. These recommendations draw on insights from the three real-world use cases discussed above and the broader cross-disciplinary analysis in the Supplementary Information on multimodal AI beyond the three use cases, providing actionable guidance for future developments.
@@
Deployment-centric development brings challenges around safety, reliability, interpretability, scalability, and ethics, particularly as multimodal AI expands in high-stake applications like healthcare and sustainability. Addressing these challenges requires implementing robust human-in-the-loop systems, developing clear standards for safety and transparency, and prioritising scalable, resource-efficient infrastructure to support increasing demands for data integration, processing, and storage.
@@
Robust data-centric development underpins deployment-centric progress by ensuring data availability, diversity, and quality. Quality multimodal data is often limited, and existing datasets often lack the diversity needed to ensure fairness in AI systems. Clear benchmarks and globally accessible datasets are needed to enhance trust, consistency, and adaptability across disciplines, facilitating reproducible research (Supplementary Fig. S1). Promoting secure data-sharing frameworks torabi2023common, open initiatives like the European Life Science Infrastructure for Biological Information crosswell2012elixir, and globally accessible multimodal data platforms will further strengthen AI’s capacity to address diverse global challenges. Knowledge graphs liang2024survey also offer promise for organising varied data formats to unify and contextualise multimodal inputs across disciplines.
@@
Model-centric development for multimodal AI faces unique challenges in effectively and efficiently fusing diverse modalities. While foundation models have proven effective for vision and language tasks, expanding them to other modalities and disciplines archit2025segment could significantly lower development barriers. Guidelines and frameworks for selecting relevant modalities, comparing multimodal versus unimodal performance, and evaluating the benefits of using many versus few modalities will help maximise model efficiency and utility.
@@
Interdisciplinary collaboration is crucial to progress, as discussed in the context of the three use cases. Standardising data integration practices, fostering cross-disciplinary knowledge exchanges, and establishing shared platforms will help bridge disciplinary divides, enabling multimodal AI to address complex societal challenges holistically.
@@
Building a dynamic and inclusive multimodal AI community fosters innovation and drives solutions to complex real-world challenges. Collaborative efforts, through regular workshops, forums, and interdisciplinary research initiatives, facilitate knowledge exchange and inspire collective problem-solving across disciplines. Engaging researchers, practitioners, and stakeholders from diverse backgrounds, particularly early-career researchers and underrepresented groups, will not only enrich the AI ecosystem but also broaden the perspectives and expertise shaping it, making it more ‘multimodal’. Community-driven initiatives Chan2020ReportingGF; de2022guidelines, including comprehensive surveys or perspective papers, provide insights that guide future developments, ensuring multimodal AI advances in responsible, impactful directions.
@@
Box 1    Strategic recommendations for advancing multimodal AI across disciplines Deployment-centric development Safety, reliability, and interpretability Challenge: Ensuring reliable, safe deployment across varied conditions and building user trust to facilitate understanding and broader adoption. Recommendation: Translate complex data and outputs into accessible formats (images, text, or speech); incorporate domain-specific knowledge; implement human-in-the-loop systems for verification and validation; conduct rigorous usability testing; develop standards for safety, reliability, and interpretability criteria; and enforce such criteria in research and peer review. Scalability and resource efficiency Challenge: Scaling multimodal AI to handle vast data volumes while optimising resource consumption, maintaining system availability, and ensuring uninterrupted communication between system components. Recommendation: Innovate scalable and resource-efficient multimodal AI solutions; promote collaboration between AI developers and hardware providers for sustainable resource use; ensure stable system availability and seamless inter-component communication; and develop cloud-based solutions and edge computing to power scalability. Ethical compliance and user preparedness Challenge: Addressing privacy, consent, and bias in applications with significant societal impacts. Recommendation: Develop clear ethical guidelines; include end-users in the design; involve human oversight and create user feedback loops for continuous solution refinement; and provide comprehensive training for reliable deployment. Data-centric development Data scarcity and access Challenge: Limited availability of high-quality multimodal datasets spanning a comprehensive range of modalities, complicated further by privacy and ethical constraints and additional heterogeneity when data sharing spans secure data environments, organisations, and regions. Recommendation: Promote open data initiatives, such as anonymised data-sharing programmes; invest in cross-discipline efforts for data collection and curation to improve availability; build data infrastructures that ensure secure, privacy-compliant, and ethics-compliant access across domains and regions; establish guidelines to standardise formats and annotations; and develop advanced tools such as knowledge graphs to integrate structured and unstructured data sources for improved contextual understanding, reasoning, and interoperability. Balanced data representation Challenge: Bias due to the limited availability of data representing underrepresented or less-studied categories (e.g. populations and regions). Recommendation: Diversify dataset development to enhance model generalisability and reduce bias; and build and enhance multimodal data platforms (e.g. UK Biobank bycroft2018uk, MIMIC johnson2016mimic, Materials Project jain2013commentary, ERA5 hersbach2020era5) to improve global data accessibility and quality. Model-centric development Multimodal fusion and modality selection Challenge: Managing diverse properties and patterns across modalities, including avoiding unnecessary modality redundancy, to maintain integrity and utility while enhancing integration. Recommendation: Assess the value of additional modalities by comparing unimodal, few-modality, and full-modality models; enhance semantic modality alignment, e.g. via large language models; develop guidelines for optimal modality selection; and address other multimodal-AI-specific modelling challenges (Fig. 2b). Foundation models (FMs) Challenge: Developing FMs beyond vision and language requires substantial resources, limiting accessibility across disciplines. Recommendation: Develop and expand multimodal FMs beyond vision and language as critical reusable infrastructure to lower barriers; and invest in downstream research and advance methods to reduce resource costs. Interdisciplinary collaboration Cross-disciplinary standards and communication Challenge: Discrepancies in multimodal data standards and communication barriers across disciplines. Recommendation: Promote cross-disciplinary standards for multimodal data integration; and organise interdisciplinary exchange events to align goals, foster collaboration, and build a thriving ecosystem. Intellectual property (IP) and workflow adaptation Challenge: IP concerns and resistance to change established workflows. Recommendation: Develop adaptive strategies and open collaborative platforms to facilitate cross-disciplinary projects.
@@
Source data for Fig. 1 and Fig. S2 is available with this paper and updated annually at https://github.com/multimodalAI/multimodal-ai-landscape.

### v2

@@ Introduction
Deploying multimodal AI systems in pandemic response translates research into actionable solutions for high-stakes environments. For instance, patient surge management can integrate biosignals, medical images, and clinical notes to optimise intensive care unit bed allocation and staff deployment. Public health decision-making can benefit from multimodal analyses of epidemiological data, EHRs, and social media trends to guide interventions such as lockdowns or resource distribution. Real-time deployment requires robust infrastructure to process asynchronous data streams and address modality-specific interruptions, such as gaps in biosignal or social media data. Adaptive algorithms and reliable systems ensure timely and accurate outbreak prediction and monitoring. Assessment through multi-centre trials and multi-disciplinary benchmarks Mincu and Roy 2022 can validate the added value of multimodal AI over unimodal systems. Measuring the accuracy of outbreak prediction or the effectiveness of intervention informed by multimodal AI can ensure trust and accountability. Further addressing scalability and privacy compliance can support sustainable and ethical deployment.
@@
Self-driving car design (Fig. 4b) exemplifies a transformative application of multimodal AI, addressing challenges in safety, efficiency, and sustainability Soni et al. 2021; Badue et al. 2021. Defining the problem involves determining how multimodal AI can substantially outperform unimodal approaches for navigating complex environments safely and efficiently. For instance, fusing sensors such as LiDAR (light detection and ranging), radar and cameras can enhance perception and decision-making in dynamic settings, especially where a single sensor may fail Yeong et al. 2021. Beyond perception, multimodal AI can support physical simulations, such as combining text, images, and videos, to model vehicle dynamics and road interactions Yang et al. 2024. In materials design, integrating structural and chemical data aids in developing lightweight, sustainable materials for vehicle construction Zhang and Xu 2022. Clearly defining these challenges ensures multimodal AI solutions are effectively tailored to the demands of autonomous driving.
@@
The deployment of multimodal AI in self-driving car design faces safety, privacy, technical, and economic constraints that must be addressed for safe and scalable adoption. Privacy concerns arise from integrating multimodal data, such as LiDAR, cameras and passenger-related information, which increases the potential for data misuse or unauthorised tracking Hansson et al. 2021. Mitigating these risks requires robust data anonymisation and adherence to privacy regulations Chowdhury et al. 2020. Technical challenges include real-time data processing in dynamic environments, such as urban intersections or adverse weather. Modality-specific failures, such as LiDAR disruptions in heavy rain, require fallback mechanisms, such as radar to maintain reliability. Self-driving systems must achieve high predictive accuracy for collision avoidance and route planning, and operate reliably across regions with varying regulations and infrastructure. Economic constraints involve balancing the cost of high-resolution sensors with scalability needs. Meeting sustainability goals requires energy-efficient operation, with trade-offs between sensor performance and power consumption. Addressing these constraints holistically enables robust and accessible deployment of self-driving technologies.
@@
In self-driving car design, task formulation maps problems and constraints to concrete AI objectives for perception, navigation, and safety. A key task is developing AI driving systems that integrate multimodal data streams, including LiDAR, radar, GPS, and cameras, for real-time navigation. These systems must dynamically process inputs to handle traffic, obstacles, and weather variability. For example, integrating road and weather data can enhance route planning and safety in adverse conditions Dey et al. 2014. Another task is predicting material properties and performance to support vehicle design. By fusing structural, experimental, and supply chain data, multimodal AI can optimise materials for lightweight, durable, and cost-effective vehicles, aligning with sustainability goals Kamran et al. 2022. Additional tasks include passenger-focused objectives, such as integrating biosignals and cabin sensors to enhance comfort and safety. Ensuring privacy compliance remains integral across all tasks.
@@
Developing multimodal AI systems for self-driving cars follows the five-step workflow above, adapted to the demands of autonomous systems. Data collection involves assembling sensor inputs (e.g. LiDAR, radar, cameras) and materials data to build comprehensive datasets. Data curation addresses challenges such as synchronising asynchronous sensor inputs and standardising materials information. Learning can employ fusion strategies to enhance system reliability, while physics-informed neural networks Raissi et al. 2019; Karniadakis et al. 2021 can integrate scientific laws to improve realistic simulations of driving scenarios and materials development. Evaluation ensures robustness through multi-metric validation across diverse scenarios, including urban environments and extreme weather conditions. Finally, interpretation in autonomous systems supports real-time explainability and transparency, enabling stakeholders to trust and audit system decisions. By aligning these steps with the unique requirements of autonomous systems, multimodal AI can ensure adaptability, safety, and performance.
@@
Real-world deployment of multimodal AI systems for self-driving cars requires live sensor integration, infrastructure readiness, and operational resilience across diverse environments. Robust in-vehicle computing and inter-vehicle communication enable reliable and coordinated performance at scale. Assessment validates multimodal AI’s added value through simulations and multi-centre trials. Testing performance under extreme conditions, such as adverse weather or high-traffic intersections, ensures reliability. Addressing scalability and privacy compliance further supports ethical, sustainable deployment, advancing societal trust in self-driving technologies Liu et al. 2019.
@@
Climate change adaptation (Fig. 4c) involves forecasting extreme weather, assessing climate risks, and managing resources to minimise environmental and socio-economic impacts Bi et al. 2023; Lam et al. 2023; Mathiesen 2018; Breitenstein et al. 2022. Defining the problem involves assessing whether multimodal AI can surpass unimodal approaches by integrating diverse data sources for deeper insights and actionable strategies. For example, multimodal AI can combine satellite imagery, historical climate records, and geospatial data to model weather patterns and predict extreme events, supporting proactive disaster management Imran et al. 2020. Integrating socio-economic data with environmental observations can enable more comprehensive climate risk assessments Tanir et al. 2024.
@@
Deploying multimodal AI for climate change adaptation must navigate privacy, technical, and economic constraints. Privacy concerns arise from integrating sensitive data, such as socio-economic records and proprietary satellite imagery. Ensuring compliance with global and regional privacy regulations and the ethical use of data supports responsible deployment. Technical challenges include the heterogeneity and sparsity of environmental and socio-economic data. Economic constraints, such as the high cost of advanced sensors, necessitate balancing low-cost and low-fidelity data sources with high-cost, high-fidelity alternatives to maintain accessibility and scalability. By addressing these constraints, multimodal AI can deliver equitable and sustainable climate solutions in diverse global contexts.
@@
For climate change adaptation, task formulation aligns complex environmental challenges with actionable AI tasks. A primary task is forecasting extreme weather events by integrating multimodal data such as satellite imagery, historical climate records, and real-time observations. These forecasts can enable timely interventions to mitigate human and economic losses Kumar et al. 2021, including early warning systems for floods that integrate geospatial, hydrological, and social data streams Tkachenko et al. 2017. Another task is predicting the socio-economic impacts of severe weather by combining spatial, market, and socio-economic data, guiding resource allocation and policy decisions Thulke et al. 2024. These tasks align multimodal AI with both technical requirements and societal priorities, advancing climate resilience and equitable adaptation strategies.
@@
Developing multimodal AI systems for climate change adaptation follows a similar five-step workflow. Data collection gathers remote sensing data, near-real-time sensor network outputs, historical climate records, and socio-economic metrics, creating comprehensive datasets for tasks such as extreme weather forecasting and early warning systems. Data curation addresses challenges such as temporal misalignment and data sparsity by standardising inputs and enriching sparse data through interpolation or integration of auxiliary sources. Multimodal learning can benefit from advanced models such as Aurora Bodnar et al. 2024, which fuse satellite imagery and meteorological data to improve atmospheric predictions. Evaluation uses multi-metric benchmarks to test system performance in diverse scenarios. For example, WeatherBench 2 provides a standard platform for assessing atmospheric prediction models Rasp et al. 2024. Interpretation ensures outputs are transparent and actionable for stakeholders, supporting evidence-based decision-making. By aligning these steps with climate-specific challenges, multimodal AI can deliver reliable, scalable, and adaptable solutions.
@@
Real-world deployment of multimodal AI for climate change adaptation translates research insights into scalable, operational systems. Early warning platforms and impact forecasting models must synchronise diverse data streams, deliver real-time outputs, and integrate with decision-making infrastructures across sectors and regions Morshed et al. 2024; Reichstein et al. 2025. Real-time monitoring requires robust infrastructure to synchronise and process diverse modalities. Deployment challenges include latency, data coverage gaps, and variable infrastructure capacity across geographies. Assessment involves stress testing and multi-centre validation to ensure reliability across environmental and socio-economic scenarios. Ethical considerations, such as prioritising vulnerable communities and ensuring equitable access to data and AI tools, help enable sustainable deployment. By overcoming deployment challenges, multimodal AI systems can empower policymakers, industries, and communities to respond proactively to climate challenges, advancing global sustainability efforts.
@@
Most existing research on multimodal AI has focused on model-centric development. While attention to data-centric development is increasing, unlocking the full potential of multimodal AI and addressing real-world challenges will ultimately require a shift towards deployment-centric development. This shift will require strategic advancements in data, model, and deployment methods, as well as concerted efforts to address multimodal-AI-specific challenges, namely modality incompleteness, multimodal heterogeneity, cross-modality alignment, modality complementarity, and multimodal privacy risk, through stakeholder engagement, interdisciplinary collaborations, and community-building. Box 1 presents strategic recommendations to advance multimodal AI across disciplines under a deployment-centric perspective. These recommendations draw on insights from the three use cases and the broader cross-disciplinary analysis in the Supplementary Information on multimodal AI beyond these three use cases, providing actionable guidance for future developments.
@@
Deployment-centric development brings challenges around safety, reliability, interpretability, scalability, and ethics, particularly as multimodal AI expands in high-stakes applications such as healthcare and sustainability. Addressing these challenges requires implementing robust human-in-the-loop systems, developing clear standards for safety and transparency, and prioritising scalable, resource-efficient infrastructure to support increasing demands for data integration, processing, and storage.
@@
Robust data-centric development underpins deployment-centric progress by ensuring data availability, diversity, and quality. High-quality multimodal data is often limited, and existing datasets often lack the diversity needed to ensure fairness in AI systems. Clear benchmarks Liang et al. 2021; de la Fuente et al. 2025 and globally accessible datasets (Fig. 5) are needed to enhance trust, consistency, and adaptability across disciplines, facilitating reproducible research. Promoting secure data-sharing frameworks Torabi et al. 2023, open initiatives such as the European Life Science Infrastructure for Biological Information Crosswell and Thornton 2012, and globally accessible multimodal data platforms will further strengthen AI’s capacity to address diverse global challenges. Knowledge graphs Liang et al. 2024a also offer promise for organising varied data formats to unify and contextualise multimodal inputs across disciplines.
@@
Model-centric development for multimodal AI faces unique challenges in effectively and efficiently fusing diverse modalities. While foundation models have proven effective for vision and language tasks, expanding them to other modalities and disciplines Archit et al. 2025 could substantially lower development barriers. Guidelines and frameworks for selecting relevant modalities, comparing multimodal versus unimodal performance, and evaluating the benefits of using many versus few modalities will help maximise model efficiency and utility. As large language models, multimodal foundation models, and generative AI systems grow in prominence Li et al. 2024; Fei et al. 2022; Narayanswamy et al. 2025; Cui et al. 2025, deployment-centric design plays an increasingly important role in guiding architectural choices, managing inference-time costs, and aligning with data and domain-specific constraints Anthropic 2024; Weisz et al. 2024. These models also increasingly rely on synthetic or weakly labelled data, or cross-modal supervision to reduce annotation costs Li et al. 2022, 2023b; Driess et al. 2023; Tan et al. 2024; Zong et al. 2024; Ding et al. 2024, reinforcing the need for deployment-aware data strategies.
@@
Stakeholder engagement and interdisciplinary collaboration are both crucial for progress, as illustrated across the three use cases. Integrating stakeholder input ensures contextual relevance and trust. Standardising data practices, fostering cross-disciplinary knowledge exchange, and developing shared platforms help bridge disciplinary gaps and enable multimodal AI to address complex societal challenges more effectively and holistically Nielsen et al. 2025; Lekadir et al. 2025.
@@
Building a dynamic and inclusive multimodal AI community fosters innovation and drives solutions to complex real-world challenges. Collaborative efforts, through regular workshops, forums, and interdisciplinary research initiatives, facilitate knowledge exchange and inspire collective problem-solving across disciplines. Engaging researchers, practitioners, and stakeholders from diverse backgrounds, particularly early-career researchers and underrepresented groups, will not only enrich the AI ecosystem but also broaden the perspectives and expertise shaping it, making it more ‘multimodal’. Community-driven initiatives Chan et al. 2020; de Hond et al. 2022, including comprehensive surveys or perspective papers, provide insights that guide future developments, ensuring multimodal AI advances in responsible, impactful directions.
@@
Box 1   Strategic recommendations for advancing multimodal AI across disciplines Deployment-centric development Safety, reliability, and interpretability Challenge: Ensuring reliable, safe deployment across varied conditions and building user trust to facilitate understanding and broader adoption. Recommendation: Translate complex data and outputs into accessible formats (images, text, or speech); incorporate domain-specific knowledge; implement human-in-the-loop systems for verification and validation; conduct rigorous usability testing; develop standards for safety, reliability, and interpretability criteria; and enforce such criteria in research and peer review Nielsen et al. 2025. Scalability and resource efficiency Challenge: Scaling multimodal AI to handle vast data volumes while optimising resource consumption, maintaining system availability, and ensuring uninterrupted communication between system components. Recommendation: Innovate scalable and resource-efficient multimodal AI solutions; promote collaboration between AI developers and hardware providers for sustainable resource use; ensure stable system availability and seamless inter-component communication; and develop cloud-based solutions and edge computing to power scalability. Ethical compliance and user preparedness Challenge: Addressing privacy, consent, and bias in applications with profound societal impacts. Recommendation: Develop clear ethical guidelines; include end-users in the design; involve human oversight and create user feedback loops for continuous solution refinement; and provide comprehensive training for reliable deployment. Data-centric development Data scarcity and access Challenge: Limited availability of high-quality multimodal datasets spanning a comprehensive range of modalities, complicated further by privacy and ethical constraints and additional heterogeneity when data sharing spans secure data environments, organisations, and regions. Recommendation: Promote open data initiatives, such as anonymised data-sharing programmes; invest in cross-discipline efforts for data collection and curation to improve availability; build data infrastructures that ensure secure, privacy-compliant, and ethics-compliant access across domains and regions; establish guidelines to standardise formats and annotations; and develop advanced tools such as knowledge graphs to integrate structured and unstructured data sources for improved contextual understanding, reasoning, and interoperability. Balanced data representation Challenge: Bias due to the limited availability of data representing underrepresented or less-studied categories (e.g. populations and regions). Recommendation: Diversify dataset development to enhance model generalisability and reduce bias; and build and enhance multimodal data platforms (e.g. UK Biobank Bycroft et al. 2018, MIMIC Johnson et al. 2016, Materials Project Jain et al. 2013, ERA5 Hersbach et al. 2020) to improve global data accessibility and quality. Stakeholder engagement and interdisciplinary collaboration Stakeholder inclusion and alignment Challenge: Limited involvement of stakeholders (e.g. domain experts, end-users, and regulators) during early-stage planning can result in solutions that are technically sound but poorly aligned with deployment contexts. Recommendation: Integrate stakeholder input across the AI lifecycle through co-design, participatory planning, and regulatory consultation to ensure contextual relevance, trust, and successful real-world adoption. Cross-disciplinary standards and communication Challenge: Discrepancies in multimodal data standards and communication barriers across disciplines. Recommendation: Promote cross-disciplinary standards for multimodal data integration; and organise interdisciplinary exchange events to align goals, foster collaboration, and build a thriving ecosystem. Intellectual property (IP) and workflow adaptation Challenge: IP concerns and resistance to change established workflows. Recommendation: Develop adaptive strategies and open collaborative platforms to facilitate cross-disciplinary projects. Model-centric development Multimodal fusion and modality selection Challenge: Managing diverse properties and patterns across modalities, including avoiding unnecessary modality redundancy, to maintain integrity and utility while enhancing integration. Recommendation: Assess the value of additional modalities by comparing unimodal, few-modality, and full-modality models; enhance semantic modality alignment, e.g. via large language models; develop guidelines for optimal modality selection aligned with deployment context and cost-performance constraints; and address other multimodal-AI-specific modelling challenges (Fig. 3b). Foundation models (FMs) Challenge: Developing FMs beyond vision and language requires substantial resources, limiting accessibility across disciplines. Recommendation: Develop and expand multimodal FMs beyond vision and language as critical reusable infrastructure to lower barriers; invest in downstream research and advance methods to reduce resource costs, e.g. using synthetic or weakly labelled data to ease training demands van Breugel et al. 2024; and pair underrepresented modalities (e.g. graphs) with more interpretable ones (e.g. language) to improve usability and broaden accessibility for non-expert stakeholders.
Figure 5: Examples of multimodal benchmark datasets. We selected two illustrative examples for each of the six disciplines to showcase diverse multimodal data and provide a starting point for multimodal AI exploration. For each dataset, we report six key attributes: year of release, data origin, data modalities, accessibility, scale of the primary modalities, and example applications. In each discipline, datasets are listed in ascending order by year of release. For the healthcare, science, and finance disciplines, we selected one small-scale dataset for quick experimentation and one larger-scale dataset for comprehensive exploration. The datasets include: TCGA (The Cancer Genome Atlas) Tomczak et al. 2015; MIMIC-IV (The Medical Information Mart for Intensive Care IV database) Johnson et al. 2023; Trafficking-10k (human trafficking advertisement dataset) Tong et al. 2017; IGDD project (Instagram data donation project) Razi et al. 2022; nuScenes (autonomous driving scene dataset) Caesar et al. 2020; DAIR-V2X (real-scenarios vehicle to everything dataset) Yu et al. 2022; Materials project (material property dataset) Jain et al. 2013; MathVista (mathematical reasoning dataset) Lu et al. 2024; iNaturalist (citizen science platform for biodiversity data) iNaturalist contributors 2025; ERA5 (European Centre for Medium-Range Weather Forecasts Reanalysis v5) Hersbach et al. 2020; Monetary policy calls (monetary policy call dataset) Mathur et al. 2022; China A-shares market (dataset of public companies listed in China A-shares market) Cheng et al. 2022. Modality definitions can vary. IGDD and ERA5 are not strictly multimodal in the traditional sense, but we treat them as such under the broader definition presented in this Perspective. The IGDD dataset contains only text data, but we consider its subtypes (chat histories and surveys) as distinct text modalities. Similarly, ERA5 comprises spatiotemporal variables such as temperature, humidity, and wind, which we treat as complementary ‘views’ offering distinct information streams for climate modelling, as adopted in prior work Chen et al. 2023; Jin et al. 2024.
@@
Source data for Fig. 1, Fig. 2, and Supplementary Fig. 1 are available with this paper and at https://github.com/multimodalAI/multimodal-ai-landscape, where they will be updated annually.